# Motion-compensated cardiac MRI using low-rank diffeomorphic flow (DMoCo)

### v2

@@ III-C Low-rank representation of velocity tensors
Here $\mathbf{p}_{n}\left(\mathbf{r}\right):\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$ are the spatial basis functions, which are independent of $s$ and $\boldsymbol{\tau}$. $R$ denotes the rank of the representation. The weights $m_{k}(\boldsymbol{\tau})$ are nonlinear functions of $\boldsymbol{\tau}$ and $s$. We represent the non-linear function using a multilayer perceptron $\mathcal{M}_{\kappa}$ with parameters $\boldsymbol{\kappa}$. The unknowns of the deformation model are $\boldsymbol{\kappa}$. Because the velocity fields are spatially smooth, we represented $\mathbf{p}_{n}(\mathbf{r})$ on a $32^{3}$ grid, which is then interpolated to the $320^{3}$ volume using trilinear interpolation.
@@ IV D-MoCo implementation
We estimate the static template $\boldsymbol{\eta}$ and the parameters of the deformation model from the measured k-space data in an unsupervised fashion. We note in (9) that the deformation model $\varphi_{\boldsymbol{\tau}};\forall\boldsymbol{\tau}$ is compactly represented in terms of low-rank factors $\mathbf{P}$ and MLP weights $\boldsymbol{\kappa}$. We introduce the fitting process in Section IV-B, which is also illustrated in Fig. 1. The motion phases are assumed to be predetermined in this step. We derive the cardiac phase $c(t)$ from the electrocardiogram (ECG). We propose to preestimate the respiratory phase from the superior-inferior (SI) k-space navigators using an autoencoder, as described below.
@@ IV-A Pre-estimation of the motion phases
We propose to estimate the respiratory phase from the SI k-space navigators using an autoencoder. In particular, we reduce the SI signal to a one-dimensional latent variable, which captures the dominant signal variation in the navigator signal that results from respiratory motion. We denote the inverse Fourier transform of multichannel SI k-space navigators, denoted by $\mathbf{x}(t)$. We filter $x(t)$ with a temporal window of 1 Hz to remove cardiac and other variations in high-frequency signals. We use a fully connected auto-encoder, with a one-dimensional latent vector $r(t)$. In particular, $r(t)$ is the one-dimensional vector that best explains the temporal variations in $x(t)$. The encoder $\mathcal{E}_{\theta_{1}}$ is an MLP with two hidden layers with 512 and 256 features, respectively. The decoder $\mathcal{F}_{\theta_{2}}$ uses the mirrored architecture, while the weights are not shared between the encoder and the decoder. The auto-encoder parameters are learned such that the loss
@@ IV-B Self-supervised Loss and Training
The unknowns of the proposed parametric deformable image model are the static template $\eta$ in (5) and the parameters of the motion model $\varphi_{\boldsymbol{\tau}}$: the spatial basis functions $\mathbf{P}$ and MLP parameters $\boldsymbol{\kappa}$. We propose to recover these parameters by the minimization of the loss $L(\eta,\mathbf{P},\boldsymbol{\kappa})$:
@@
The first term is the data consistency between the deformed templates and the k-space data. The second term is a penalty to encourage the velocity fields to be path independent as described in Section III-B, and the third is a spatial smoothness penalty on the static template. Here, $\varphi_{\boldsymbol{\tau}}$ is the deformation obtained by integrating the velocity field along a straight line-segment $\mathcal{C}_{\tau}$ as in (7). $\psi_{\boldsymbol{\tau}}$ is the integral along a perturbed path $\mathcal{D}_{\boldsymbol{\tau}}$ as in (8). The second term involves an expectation over all possible paths between 00 and $\boldsymbol{\tau}$. We evaluate both integrals using Euler’s summation. The random path is obtained by perturbing the samples on the line by Gaussian noise of variance 10 percent of the step size between the time points. We use different random paths at each iteration of the optimization to encourage the deformation to be independent of the paths.
We minimize the above loss using stochastic gradient descent. At each iteration, we pick a random $\boldsymbol{\tau}\in[0,1]^{2}$. We choose $K$ readout groups whose phase is the closest to $\tau$ in the Euclidean sense to define the loss and hence the gradient. In this work, we chose $K=20$. The hyperparameters are chosen by trial and error on one of the datasets and were kept fixed for the other datasets. We observe that the results were not very sensitive to $\lambda_{1}$, provided that it is large enough to ensure that the deformations are independent of the path. The parameter $\lambda_{2}$ was chosen to yield a good compromise between streak artifacts in the reconstructed template $\eta$ and blurring.
@@ IV-C Initialization of motion fields for fast convergence
To speed up the training, we introduce a fast approach to initialize the velocity fields. In particular, we group the k-space readouts to five cardiac and three respiratory motion bins, in an approach similar to XD-GRASP [11]. We reconstructed the images corresponding to each of the bins using a smoothness regularized reconstruction:
$\boldsymbol{\rho}_{\boldsymbol{\tau}_{k}}=\arg\min_{\mathbf{\rho}}\left\|A_{\boldsymbol{\tau}_{k}}\left(\boldsymbol{\rho}_{\boldsymbol{\tau}_{k}}\right)-\mathbf{b}_{\tau_{k}}\right|_{2}^{2}+\lambda_{3}\left\|\nabla\boldsymbol{\rho}_{\mathbf{\tau}_{k}}\right\|_{2}^{2}$ (12)
We note that each of the bins are heavily undersampled. We hence choose a high enough $\lambda_{3}$ to obtain a low-resolution reconstruction, free of visible alias artifacts. We set $\eta=\boldsymbol{\rho}_{0}$ and solve for $\boldsymbol{\tau}$ and $\boldsymbol{\kappa}$ by minimizing the loss:
@@
Because this optimization problem does not involve multichannel NUFFT steps, this initialization strategy is computationally much more efficient than that of (11). We use stochastic gradient descent to derive $\mathbf{P}$ and $\mathbf{\kappa}_{\rm init}$. We note that these parameters are used only to initialize the optimization scheme.
Figure 1: DMoCo architecture: (a) The static template volume $\boldsymbol{\eta}$. (b) The velocity fields at a specific phase $\boldsymbol{\tau}$ is modeled as the product of the static basis vectors $\mathbf{P}$ and its weights $m_{n}(s,\boldsymbol{\tau})$. The weights are derived from the cardiac and respiratory phase $\boldsymbol{\tau}=(c,r)$ using a multilayer perceptron with parameters $\boldsymbol{\kappa}$. (c) The deformation field are modeled as a forward diffeomorphic flow, denoted by (7) along the straightline path $\mathcal{C}_{\boldsymbol{\tau}}$. (d) The static image volume $\boldsymbol{\eta}$ is interpolated to each phase using trilinear interpolation. (e) The deformation is also derived along an alternate stochastic path $\mathcal{D}_{\boldsymbol{\tau}}$; the difference between the two different deformation estimates at $\boldsymbol{\tau}$ are used as a penalty, which is the second term in (11). This term ensures that the velocity integrals are path independent, which further constraints the deformation estimates. (f) The forward operator involves multiplication by the coil sensitivity maps and NUFFT. Here, we consider $K$ readout groups with the closest phase to $\boldsymbol{\tau}$ and evaluate the $\ell_{2}$ error from the corresponding measurements measured k-space values. A smoothness penalty is also added to the static template.
@@ IV-D Image Acquisition
The data are acquired using a 3D radial balanced steady-state free precession (bSSFP) research acquisition scheme using a spiral phyllotaxis view ordering on a 1.5T MRI scanner (MAGNETOM Aera, Siemens Healthcare, Forchheim, Germany) system using a 32-channel cardiac array. The view order consisted of a superior-inferior (SI) k-space navigator, acquired in every 22 spokes. The temporal resolution of each group of 22 spokes is approximately   70 $\mathrm{ms}$. The number of radial readouts acquired per spoke were 5749, which corresponds to a fixed acquisition time of 5:58 minutes. The main acquisition parameters were: TR/TE of 3.14/1.6 $\mathrm{ms}$, radio frequency excitation angle of $60^{\circ}$, resolution of 1.5 $\mathrm{mm}^{3}$ isotropic, field of view (FOV) of 240 $\mathrm{mm}^{3}$, and readout bandwidth of 840 Hz/pixel [14, 25].
@@ IV-E State of the art (SOTA) methods used for comparison
We compare our method against MoCo-SToRM [34]. In MoCo-SToRM, each volume in the time series is modeled as the deformation of a static template, similar to the proposed approach. The main distinction is that the deformation fields are directly represented by a compact model in MoCo-STORM, unlike the low-rank representation of the velocity fields that are integrated to obtain the deformation fields. To make the comparison fair, we represented the deformations in MoCo-STORM using a low-rank representation
@@
similar to (9). Here, $\mathbf{Q}(\mathbf{r})$ are the $R$-dimensional spatial basis vectors and $\mathcal{N}_{\kappa}$ is an MLP with parameters $\kappa$.
We also implemented a motion-resolved (MoR) approach similar to [11]. The cardiac data is binned to three respiratory phases and five cardiac phases [11]. With a short acquisition of 5:58 minutes, it is challenging to resolve more than five cardiac phases. The $\hat{\boldsymbol{\rho}}_{\boldsymbol{\tau}_{ij}}$ volumes are recovered using (15).
@@
Here we use the same notation as the above methods. However, we use a temporal regularization between the bins $\left\|\nabla_{ij}\boldsymbol{\rho}\right\|_{1}$ denotes a total variation penalty. The above motion resolved reconstruction was done using a conjugate gradient algorithm. We also denote the variables as with subscript $\tau_{ij}$ to show that the reconstruction is discretized to a fixed set of bins.
@@ V Experimental setup
A challenge in quantitatively validating the proposed approach in an in vivo setting is the lack of fully sampled cardiac and respiratory resolved 3D datasets; the acquisition of these data is impossible. Such data sets are not available to the best of our knowledge. We therefore restrict quantitative comparisons to a numerical phantom generated using XCAT simulations [1] and a 2D annulus with cardiac and respiratory motion Fig. 2. We also compare the proposed method with state-of-the-art methods on in vivo datasets acquired from patients, where we use their breath-held 2D cine acquisitions as reference.
Figure 2: Low-rank nature of the velocity fields. We show the estimated velocity and deformation maps for a 2D annulus numerical phantom (top-left) and the 3D in-vivo dataset (right). The top rows correspond to the images in three different phases, with the right-most image being the template. The middle rows correspond to the estimated deformations, overlaid on the template, while the bottom rows show the velocities in these phases. We note that the velocity fields (bottom rows) are almost the same across different phases, while the corresponding deformations are significantly more complex and changes drastically. In particular, the integration of a relatively static velocity field as in (7) can account for complex time-varying deformation fields. We performed a low-rank approximation of the velocity fields (orange), deformation fields (blue), and the image time series (green) and evaluated the corresponding errors. We observe that a significantly lower rank is sufficient to yield smaller errors, compared to the image time series and even the deformation fields. Here, a rank of 3 on the velocity tensor gave better results than a rank of 10 on the deformation fields. These results show that a deformation model using low-rank modeling of the velocities is very efficient in representing the image time series.
@@
An XCAT phantom was used to generate 5749 volumes of repeating respiratory and cardiac patterns. We assumed a respiratory period of 5s and a heart rate of 60 beats per minute. We simulated the k-space measurements by multiplying the phantom by $B_{1}$ profile and simulated coil sensitivities of a birdcage coil sensitivity mask with 6 coils and computing the nonuniform fast Fourier transformation of each radial group. Noise of standard deviation $\sigma=5\times 10^{-3}$ was added to the k-space measurements. The image quality metrics were then calculated against the ground truth as well as the respiratory displacements (RD). RD was assessed by measuring the displacement of the left diaphragm dome for the fitted volumes. In addition, we compared the effect of modeling the flow and velocity fields. The total simulated scan time was about six minutes and used the same FOV and resolution as in the in vivo study.
@@
We use a 2D annulus phantom, considered in Fig. 2, to compare the low-rank approximations of the velocity fields, deformation fields, and the image time series. The 2D setting makes the computational complexity and memory demand of the low-rank approximations manageable. Here, we consider an annulus that is expanding and contracting to mimic cardiac motion with 60 beats per minute. The cardiac motion was generated using two circles that contract at different rates. The outer circle’s radius decreases by 10% of the FOV while the radius of the inner circle decreases by 14% of the FOV. The phantom also consists of a circular section at the bottom to mimic the liver. We also consider contracting motion with a period of 4s and a displacement of 24%

### v3

@@ III-D Low-rank representation of velocity tensors
Here $\mathbf{p}_{n}\left(\mathbf{r}\right):\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$ are the spatial basis functions, which are independent of $s$ and $\boldsymbol{\tau}$. $R$ denotes the rank of the representation. The weights $m_{k}(\boldsymbol{\tau})$ are nonlinear functions of $\boldsymbol{\tau}$ and $s$. We represent the non-linear function using a multilayer perceptron $\mathcal{M}_{\kappa}$ with parameters $\boldsymbol{\kappa}$. The unknowns of the deformation model are $\boldsymbol{\kappa}$. Because the velocity tensors are spatially smooth, we represented $\mathbf{p}_{n}(\mathbf{r})$ on a $32^{3}$ grid, which is then interpolated to the $320^{3}$ volume using trilinear interpolation.
@@ IV D-MoCo implementation
We estimate the static template $\boldsymbol{\eta}$ and the parameters of the deformation model from the measured k-space data in an unsupervised fashion. The preestimation of the motion phases is described in Section IV-A. We formulate image recovery as an optimization scheme as described in IV-B, where we estimate unknowns of the model: (a) low-rank factors $\mathbf{P}$ and MLP weights $\boldsymbol{\kappa}$ of the deformation model in (9) that the deformation model, and (b) the image template $\eta$. The fitting process is illustrated in Fig. 1.
@@ IV-A Pre-estimation of the motion phases
We estimate the respiratory phase from the SI k-space navigators using an auto-encoder. In particular, we reduce the SI signal to a one-dimensional latent variable, which captures the dominant signal variation in the navigator signal that results from respiratory motion. We denote the inverse Fourier transform of multichannel SI k-space navigators, denoted by $\mathbf{x}(t)$. We filter $x(t)$ with a temporal window of 1 Hz to remove cardiac and other variations in high-frequency signals. We use a fully connected auto-encoder, with a one-dimensional latent vector $r(t)$. In particular, $r(t)$ is the one-dimensional vector that best explains the temporal variations in $x(t)$. The encoder $\mathcal{E}_{\theta_{1}}$ is an MLP with two hidden layers with 512 and 256 features, respectively. The decoder $\mathcal{F}_{\theta_{2}}$ uses the mirrored architecture, while the weights are not shared between the encoder and the decoder. The auto-encoder parameters are learned such that the loss
@@
The MLP $\mathcal{M}_{\boldsymbol{\kappa}}$ in Section III-D uses an input layer, which transforms the phase vector $\boldsymbol{\tau}=[c(t),r(t)]^{T}$ to the 3D vector $[\sin(2\pi c(t)/C),\cos(2\pi c(t)/C),r/R]$, where $R$ is the maximum magnitude of $r(t)$. This transformation preserves the complex harmonic nature of cardiac motion and the beat-to-beat variability of real-time imaging.
@@ IV-B Self-supervised Loss and Training
The unknowns of the proposed parametric deformable image model are the static template $\eta$ in (5) and the parameters of the motion model $\varphi_{\boldsymbol{\tau}}$: the spatial basis functions $\mathbf{P}$ and the MLP parameters $\boldsymbol{\kappa}$. We propose to recover these parameters by the minimization of the loss $L(\eta,\mathbf{P},\boldsymbol{\kappa})$:
@@
The first term is the data consistency between the deformed version $\eta(\varphi_{\boldsymbol{\tau}}(\mathbf{r}))$ of the template $\eta(\mathbf{r})$ at the phase $\boldsymbol{\tau}$ and the k-space data $\mathbf{b}_{\boldsymbol{\tau}}$ at that phase, and the third is a spatial smoothness penalty on the static template. The second term in the above equation is a penalty to encourage the velocity tensors to be path-independent as described in Section III-C. Here, $\varphi_{\boldsymbol{\tau}}$ is the deformation obtained by integrating the velocity tensor along a straight-line segment $\mathcal{C}_{\tau}$ as in (7). $\psi_{\boldsymbol{\tau}}$ is the integral along a perturbed path $\mathcal{D}_{\boldsymbol{\tau}}$ as in (8). The second term involves an expectation of all possible paths between 00 and $\boldsymbol{\tau}$. We evaluated both integrals using Euler’s summation. The random path is obtained by perturbing the samples on the line by Gaussian noise of variance 10 percent of the step size between the time points. We use different random paths at each iteration/epoch of the optimization to encourage the deformation to be independent of the paths.
We minimize the above loss using stochastic gradient descent. In each iteration, we pick a random $\boldsymbol{\tau}\in[0,1]^{2}$. We choose $K$ readout groups whose phases are the closest to $\tau$ in the Euclidean sense to define the loss and hence the gradient. In this work, we chose $K=20$. The hyperparameters are chosen by trial and error on one of the datasets and are kept fixed for the other datasets. We observe that the results were not very sensitive to $\lambda_{1}$, provided that it is large enough to ensure that the deformations are independent of the path. The parameter $\lambda_{2}$ was chosen to yield a good compromise between streak artifacts in the reconstructed template $\eta$ and blurring.
@@ IV-C Initialization of motion fields for fast convergence
To speed up the training, we introduce a fast approach to initialize the velocity tensors. In particular, we group the k-space readouts to five cardiac and three respiratory motion bins, in an approach similar to XD-GRASP [30]. We reconstructed the images corresponding to each of the bins using a smoothness regularized reconstruction:
$\boldsymbol{\rho}=\arg\min_{\mathbf{\rho}}\sum_{k}\left\|A_{\boldsymbol{\tau}_{k}}\left(\boldsymbol{\rho}_{\boldsymbol{\tau}_{k}}\right)-\mathbf{b}_{\tau_{k}}\right|_{2}^{2}+\lambda_{3}\left\|\nabla\boldsymbol{\rho}\right\|_{2}^{2}$ (12)
We note that each of the bins are heavily undersampled. We therefore chose a high value of $\lambda_{3}$ to obtain blurred images that are free of visible alias artifacts. We note that these images are only used to estimate the initial velocity tensors. We set the template as the image at $\boldsymbol{\tau}=0$ ($\eta=\boldsymbol{\rho}_{0}$) and solve for $\boldsymbol{\tau}$ and $\boldsymbol{\kappa}$ by minimizing the loss:
@@
Because this optimization problem does not involve multichannel NUFFT, this initialization strategy is computationally much more efficient than that of (11). We use stochastic gradient descent to derive $\mathbf{P}$ and $\mathbf{\kappa}_{\rm init}$. We note that these parameters are used only to initialize the optimization scheme.
Figure 1: DMoCo architecture: (a) The static template volume $\boldsymbol{\eta}$. (b) The velocity tensors at a specific phase $\boldsymbol{\tau}$ is modeled as the product of the static basis vectors $\mathbf{P}$ and its weights $m_{n}(s,\boldsymbol{\tau})$. The weights are derived from the cardiac and respiratory phase $\boldsymbol{\tau}=(c,r)$ using a multilayer perceptron with parameters $\boldsymbol{\kappa}$. (c) The deformation field are modeled as a forward diffeomorphic flow, denoted by (7) along the straightline path $\mathcal{C}_{\boldsymbol{\tau}}$. (d) The static image volume $\boldsymbol{\eta}$ is interpolated to each phase using trilinear interpolation. (e) The deformation is also derived along an alternate stochastic path $\mathcal{D}_{\boldsymbol{\tau}}$; the difference between the two different deformation estimates at $\boldsymbol{\tau}$ are used as a penalty, which is the second term in (11). This term ensures that the velocity integrals are path independent, which further constraints the deformation estimates. (f) The forward operator involves multiplication by the coil sensitivity maps and NUFFT. Here, we consider $K$ readout groups with the closest phase to $\boldsymbol{\tau}$ and evaluate the $\ell_{2}$ error from the corresponding measurements measured k-space values. A smoothness penalty is also added to the static template.
@@ IV-D Image Acquisition
[…]
@@ IV-E State of the art (SOTA) methods used for comparison
We compare our method against MoCo-SToRM [23]. In MoCo-SToRM, each volume in the time series is modeled as the deformation of a static template, similar to the proposed approach. The main distinction is that the deformation fields are directly represented by a compact model in MoCo-STORM, unlike the low-rank representation of the velocity tensors that are integrated to obtain the deformation fields. To make the comparison fair, we represented the deformations in MoCo-STORM using a low-rank representation
@@
similar to (9). Here, $\mathbf{Q}(\mathbf{r})$ are the spatial basis vectors of dimension $R$ and $\mathcal{N}_{\kappa}$ is an MLP with parameters $\kappa$.
We also implemented a motion-resolved approach (MoR) that is conceptually similar to [30]. The cardiac data are grouped into three respiratory phases and five cardiac phases [30]. We note that with a short acquisition of 5:58 minutes, it is challenging to resolve more than five cardiac phases. The $\hat{\boldsymbol{\rho}}_{\boldsymbol{\tau}_{ij}}$ volumes are recovered using (15).
@@
Here we use the same notation as the above methods. However, we use a temporal regularization between the bins $\left\|\nabla_{ij}\boldsymbol{\rho}\right\|_{1}$ denotes a total variation penalty. The above motion-resolved reconstruction was done using a conjugate gradient algorithm. We also denote the variables as with the subscript $\tau_{ij}$ to show that the reconstruction is discretized to a fixed set of bins.
@@ V Experimental setup
A challenge in quantitatively validating the proposed approach in an in-vivo setting is the lack of fully sampled cardiac and respiratory resolved 3D datasets; the acquisition of these data is impossible. We therefore restrict quantitative comparisons to a numerical phantom generated using XCAT simulations [32] and a 2D annulus with cardiac and respiratory motion Fig. 2. We also compare the proposed method with state-of-the-art methods using patient data sets. We note that quantitative image comparisons are challenging in this case due to drastically different tissue contrasts and mismatches in alignment. We instead visually compare 2D slices extracted from the reconstructed volumes that closely match with breath-held 2D cine acquisitions. We also performed segmentations of the left ventricles and perform a preliminary comparison with similar measures from 2D cine.
Figure 2: Impact of low-rank approximation of velocity tensors, deformation fields, and image series: We show the estimated velocity and deformation maps for a 2D annulus numerical phantom (top-left) and the 3D in-vivo dataset (right). The top rows correspond to the images in three different phases, with the right-most image being the template. The middle rows correspond to the estimated deformations (shown by arrows), overlaid on the template, while the bottom rows show the velocities in these phases. We note that the velocity tensors (bottom rows) are almost the same across different phases, while the corresponding deformations are significantly more complex and changes drastically. In particular, the integration of a relatively static velocity tensor as in (7) can account for complex time-varying deformation fields. In the plot (bottom-left) we performed a low-rank approximations with differing rank of the velocity tensors, deformation fields, and the image time series from the annulus phantom. We computed the corresponding reconstructions, which were compared against the original images. The plots show the errors corresponding to low-rank approximations on the velocity tensors in orange, deformation fields in blue, and images in green. We observe that a significantly lower rank on the velocity tensors is sufficient to yield smaller errors, compared to the image time series and even the deformation fields. Here, a rank of 3 on the velocity tensor gave better results than a rank of 10 on the deformation fields. These results show that a deformation model using low-rank modeling of velocities is very efficient in representing the image time series.
@@
We generated numerical data using the XCAT framework [32], which consists of 5749 volumes, with respiratory and cardiac patterns. We assumed a respiratory period of 5 s and a heart rate of 60 beats per minute. We simulated the k-space measurements by multiplying the phantom by $B_{1}$ profile and simulated coil sensitivities of a birdcage coil sensitivity mask with 6 coils, followed by a non-uniform fast Fourier transformation (NUFFT) of each radial group. Noise of standard deviation $\sigma=5\times 10^{-3}$ was added to the k-space measurements. The image quality metrics were then calculated against the ground truth as well as the respiratory displacements (RD). RD was assessed by measuring the displacement of the left diaphragm dome for the fitted volumes. In addition, we compared the effect of modeling the flow and velocity tensors. The total simulated scan time was about six minutes and used the same FOV and resolution as in the in vivo study.
@@
We use a 2D annulus phantom, considered in Fig. 2, to compare the low-rank approximations of the velocity tensors, deformation fields, and the image time series. The 2D setting makes the computational complexity and memory demand of the low-rank approximations manageable. Here, we consider an annulus that is expanding and contracting to mimic cardiac motion with 60 beats per minute. The cardiac motion was generated using two circles that contract at different rates. The radius of the outer circle decreases by 10% of the FOV while the radius of the inner circle decreases by 14% of the FOV. The phantom also consists of a circular section at the bottom to mimic the liver. We also consider a contracting motion with a period of 4s and a displacement of 24%